OpenStudy (anonymous):

A building has a ramp to its front doors to accommodate the handicapped. If the distance from the building to the end of the ramp is 13 feet and the height from the ground to the front doors is 5 feet, how long is the ramp? (Round to the nearest tenth.)

OpenStudy (anonymous):

You can use the Pythagorean Theorem. a^2 + b^2 = c^2 13^2 + 5^2 = c^2

OpenStudy (anonymous):

So, \[x = \sqrt{13^{2} + 5^{2}}\]

OpenStudy (anonymous):

sqrt194 = 13.93

OpenStudy (anonymous):

You got it! Good job!

OpenStudy (anonymous):

oh thats it? dang... thanks so much lol. im dumb sometimes

OpenStudy (anonymous):

No, we just forget sometimes. The main thing is that you know.

OpenStudy (anonymous):

Well, it looks like you're all set here, so, good working with you and good luck to you in all your future studies! @kjuchiha